# Each generic polytope in $\mathbb{R}^{3}$ has a point with ten normals to the boundary

### v3

@@ Abstract.
We treat the same problem for convex polytopes in $\mathbb{R}^{3}$ and prove that each generic polytope has a point in its interior with $10$ normals to the boundary.
@@ 1. Introduction
In this paper we solve a similar problem for compact convex polytopes with non empty interior.
@@
Let $\mathbf{P}\subset\mathbb{R}^{3}$ be a compact convex polytope with non empty interior, let $y\in Int\leavevmode\nobreak\ \mathbf{P}$. Let a support plane $h$ to $\mathbf{P}$ and a point $z\in h\cap\mathbf{P}$ be such that the line $yz$ is orthogonal to $h$. Then $yz$ is called a normal to the boundary of $\mathbf{P}$ emanating from $y$. The point $z$ is called the base of the normal.
@@ 2.1. Active regions, bifurcation set and its sheets
The bifurcation set $\mathcal{B}(\mathbf{P})$ is the union of the boundaries of active regions. It is the PL counterpart of the focal set that appears in the smooth case.
@@
    Local minima are attained at some faces of $\mathbf{P}$.
@@
    Saddles are attained at some edges of $\mathbf{P}$.
@@
We say that a point $y\in Int(\mathbf{P})$ projects to a face $F$ (to an edge $e$, respectively) if the orthogonal projection of $y$ to the affine hull of $F$ (of $e$, respectively) lies in the interior of $F$ (of $e$).
@@
    A point $y\in\mathbf{P}$ projects to the interior of a face iff the face contributes a minimum of $SQD_{y}$.
@@
In the paper we restrict ourselves by generic polytopes:
@@
Important observations are: (1) For a generic polytope, the sheets of $\mathcal{B}(\mathbf{P})$ intersect transversally. (2) Any polytope can be turned to a generic one by a close to $id$ projective transform. (3) A sheet of $\mathcal{B}$ related to a face contains no vertices of $\mathbf{P}$ except for vertices of the face.
@@ 2.2. Associated spherical geometry
Let $V$ be a vertex of $\mathbf{P}$. Let a point $y$ lie in the interior of $\mathbf{P}$, close to the vertex $V$. Here and in the sequel we denote by $Y$ the intersection of the small sphere centered at $V$ and the ray $Vy$.
@@
The core $Core(P)$of a convex spherical polygon $P$ is the set of its interior points $Y$ such that the distance from $Y$ to each point from $\partial P$ does not exceed $\pi/2$.
@@
    $V$ is a local maximum of $SQD_{y}$ iff $Y\in Core(Spher(V))$.
@@
    A point $Y\in Int(P)$ has a short maximum (respectively, short minimum) at $X\in\partial P$ if $X$ is a local maximum (respectively, a local minimum) of $SQD^{sph}_{Y}$, and $|YX|<\pi/2$.
@@
    A vertex $V$ of polytope $\mathbf{P}\subset\mathbb{R}^{3}$ is skew if the associated spherical polygon $Spher(V)$ is skew.
@@
For a fixed $Y$, maxima and minima of $SQD_{Y}^{sph}$ alternate on $\partial P$. We say that a minimum and a maximum are neighbours if there is no other maxima and minima on a connecting arch of $\partial P$.
Example. Skew triangles were studied in [5] in details. A skew triangle is depicted in Fig 2.
@@ 3. Proof of Theorem 1
    Therefore, for every point on the ray $VY$, the vertex $V$ is a local maximum. Take a point $y$ lying on the ray very close to $V$ and let it travel along the ray. We may assume that $y$ never crosses two sheets of the bifurcation set at a time. If this is not the case, perturb $y$ a little bit.
    Initially $SQD_{y}$ has three minima $m_{1}$, $m_{2}$, and $m_{3}$, lying on the faces incident to $V$ , three saddles $s_{1}$, $s_{2},s_{3}$, lying on edges that correspond to $A_{1},A_{2},A_{3}$, and one local maximum $V$. Since $V$ is not the global maximum, there is at least one other maximum. Altogether there are at least 8 critical points of $SQD_{y}$, or, equivalently, 8 normals emanating from $y$.
@@
    (c) One of the saddles $s_{i}$ dies. This happens in two cases: (i) $s_{i}$ meets a (local) maximum of $SQD_{y}$. The maximum cannot be $V$ since $s_{i}$ slides away from $V$. It is not any other maximum $V^{\prime}$. Indeed, recall that $s_{i}$ is the orthogonal projection of $y$ to the edge. If $s_{i}$ meets $V^{\prime}$ before $y$ leaves $P$, then $|Vy|>|VV^{\prime}|$. Then $V^{\prime}$ can’t be maximum of $SQD_{y}$ when $y=V$. (ii) $s_{i}$ meets a minimum. It can be none of $m_{1}$, $m_{2}$, and $m_{3}$, so right before the bifurcation, the function $SQD_{y}$ has $4$ minima and $2$ maxima, which proves the claim.
@@
    Each skew polygon has exactly two acute angles, say, at vertices $M_{1}$ and $M_{2}$,
@@
By Proposition 2, (1) each vertex of $\mathbf{P}$ is incident to exactly two acute edges. Therefore, acute edges form a closed broken line on the boundary of $\mathbf{P}$. Denote the broken line by $\mathcal{A}$.
@@
    $SQD_{y}:\partial P\rightarrow\mathbb{R}.$
@@
(2) One way is clear: a maximum of $SQD_{y}$ is also a maximum of $SQD_{y,\mathcal{A}}$. The converse follows from Proposition 2, (3) and Lemma 5.
@@
There exists a binormal $VW$ of $P$ such that $V$ is a vertex, and $W$ lies in the interior of some edge.
@@
Bimormals of $\mathbf{P}$ correspond to critical points of the width function of $\mathbf{P}$.  [^2] The width function is clearly not a smooth function. So our aguments rely on the concept of Clarke subdifferential borrowed from non-smooth analysis, see e.g. [1], which contains also the definition of critical point and the regular interval theorem.
@@
(1) If the endpoints of a binormal are two vertices of $\mathbf{P}$, then the binormal realises a local maximum of the width function.
(2) If the endpoints of a binormal are a vertex of $\mathbf{P}$ and a point in the interior of a face, then the binormal realises a local minimum.
(3) If the endpoints of a binormal lie in the interiors of some (non-parallel, because of genericity) edges of $K$, then the binormal also realises a local minimum.
@@
Observe that $VW$ belongs to the active region $\mathcal{AR}(V)$, that is, for every point on the binormal, $V$ is a local maximum.
@@
Take a point $y$ lying on the binormal very close to the vertex $V$. By Lemmata 5 and 3, $SQD_{y}$ has two saddles $s_{1},s_{2}$ on the acute edges incident to $V$, a saddle $s_{3}=W$, two minima $m_{1},m_{2}$ on the faces incident to $V$, and a local maximum at $V$. Since $V$ is not a global maximum, there is some other global maximum; altogether there are at least $8$ normals emanating from $y$.
Figure 3. Binormal $VW$
@@
Let $VV^{\prime}$ and $VV^{\prime\prime}$ be two acute edges incident to $V$. For a point $y$ lying close to $V$ none of $V^{\prime}$ and $V^{\prime\prime}$ is a (local) maximum of $SQD_{y}$.
@@
Keep moving until the first event, which is either a bifurcation (that is, a crossing of a sheet of $\mathcal{B}(P)$), or the point $y$ reaches $W$. We have the following case analysis.
@@
      One of the saddles $s_{i}$ meets a local maximum, and the two critical points die. This saddle cannot be $s_{3}$ since $s_{3}$ does not move. It can be neither $s_{1}$ nor $s_{2}$ by Lemma 8.
@@
One repeats the arguments from Case 1 for a point $y$ travelling along the binormal shifted a little bit inwards the polytope. The same bifurcation arguments survive: if $y$ crosses a sheet of $\mathcal{B}(P)$ on its way, we get $10$ emanating normals.
So the only remaining case is the one when there are no bifurcations on the (shifted) binormal. Here we need additional arguments since one of the two new minima coincides with either $m_{1}$ or $m_{2}$. However we conclude that for a point $y$ on the (shifted) binormal we have (at least) $3$ minima, $2$ maxima, and $3$ saddles. Moreover, these $3$ saddles are attained on acute edges of $P$. That is, we have three minima of $SQD_{y,\mathcal{A}}$, see Lemma 6. Since $\mathcal{A}$ is a topological circle (or a disjoint union of topological circles), there are necessarily three maxima of $SQD_{y,\mathcal{A}}$. By Lemma 6, they are maxima of $SQD_{y}$ as well.
@@ 4. Proof of Proposition 2
A bigon is a spherical polygon with two edges. It is bounded by two halves of big circles. The minimal bigon $Big(P)$ of a spherical polygon $P$ is the bigon containing $P$ which is minimal by inclusion, or, equivalently, with the minimal angle between its edges.
@@
Let $A$ be a vertex of $P$. Set $\mathcal{AR}^{spher}(A)$ be the set of points $Y\subset P$ such that $A$ is a short local maximum of $SQD_{Y}^{spher}$.
@@
It remains to analyze triangled and quadrilaterals.
@@
Assume that $A$ lies in the same half of bigon as $M_{1}$. Then we have:$|M_{1}p|<\pi/2$, $|M_{1}A|\leqslant\pi/2$, $\angle AM_{1}p<\pi/2$, which implies $|Ap|<\pi/2$. ∎

### v4

@@ Abstract.
We treat an analogous problem for convex polytopes in $\mathbb{R}^{3}$ and prove that each generic polytope has a point in its interior with at least $10$ normals to the boundary.
@@
The proof is based on piecewise linear analog of Morse theory, analysis of bifurcations, and some combinatorial tricks.
@@ 1. Introduction
In this paper we solve a similar problem for compact convex polytopes with non-empty interior.
@@
Let $\mathbf{P}\subset\mathbb{R}^{3}$ be a compact convex polytope with non-empty interior, let $y\in Int\leavevmode\nobreak\ \mathbf{P}$. Let h be a support plane to $\mathbf{P}$ and a point $z\in h\cap\mathbf{P}$ be such that the line $yz$ is orthogonal to $h$. Then $yz$ is called a normal to the boundary of $\mathbf{P}$ emanating from $y$. The point $z$ is called the base of the normal.
@@
The main idea in a nutshell looks like this: (1) take a very special point $y$ in the polytope $\mathbf{P}$ with some guaranteed number of emanating normals. (2) Force $y$ to move along a carefully chosen straightline trajectory. At bifurcation moments od the movement, the emanating normals born or die. Analyze this birth-death process.
We use this idea twice: in Proposition 1, where $y$ travels along a special ray emanating from a vertex of $\mathbf{P}$, and in the proof of the main theorem, where $y$ travels along a special (saddle-type) binormal of $\mathbf{P}$.
Finally, a combinatorial trick puts the pieces together.
@@ 2.1. Active regions, bifurcation set and its sheets
The bifurcation set $\mathcal{B}(\mathbf{P})$ is the union of the boundaries of active regions. It is the piecewise linear counterpart of the focal set that appears in the smooth case.
@@
    Local minima of $SQD_{y}$ are attained at some faces of $\mathbf{P}$.
@@
    Saddles of $SQD_{y}$ are attained at some edges of $\mathbf{P}$.
Convention: from now on, we call local maxima (local minima) just maxima and minima for short.
@@
We say that a point $y\in Int(\mathbf{P})$ projects to a face $F$ (to an edge $e$, respectively) if the orthogonal projection of $y$ to the affine hull of $F$ (of $e$, respectively) lies in the interior of the face $F$ (or of the edge $e$).
@@
    A point $y\in\mathbf{P}$ projects to a face iff the face contributes a minimum of $SQD_{y}$.
@@
In the paper we restrict ourselves to generic polytopes:
@@
Important observations are: (1) For a generic polytope, the sheets of $\mathcal{B}(\mathbf{P})$ intersect transversally. (2) Any polytope can be turned to a generic one by a close to $id$ projective transform. (3) If $\mathbf{P}$ is generic, then any sheet of $\mathcal{B}(\mathbf{P})$ related to a face contains no vertices of $\mathbf{P}$ except for vertices of the face.
@@ 2.2. Associated spherical geometry
Let $V$ be a vertex of a polytope $\mathbf{P}$. Let a point $y$ lie in the interior of $\mathbf{P}$, close to the vertex $V$. Here and in the sequel we denote by $Y$ the intersection of the small sphere centered at $V$ and the ray $Vy$.
@@
The core of a convex spherical polygon $P$ is defined as follows:
[…]
@@
    $V$ is a maximum of $SQD_{y}$ iff
    $Y\in Core(Spher(V))$.
@@
    A point $Y\in Int(P)$ has a short maximum (respectively, short minimum) at $X\in\partial P$ if $X$ is a maximum (respectively, a minimum) of $SQD^{sph}_{Y}$, and $|YX|<\pi/2$.
@@
    A vertex $V$ of a polytope $\mathbf{P}\subset\mathbb{R}^{3}$ is skew if the associated spherical polygon $Spher(V)$ is skew. Otherwise $V$ is called nice.
@@
For a fixed $Y$, maxima and minima of $SQD_{Y}^{sph}$ alternate on $\partial P$. We say that a minimum and a maximum are neighbours if there is no other maxima and minima on a connecting arc of $\partial P$.
Example. Skew triangles were studied in [5] in detail. A skew triangle is depicted in Fig 2.
@@ 3. Proof of Theorem 1
    Therefore, for every point on the ray $VY$, the vertex $V$ is a maximum. Take a point $y$ lying on the ray very close to $V$ and let it travel along the ray. We may assume that $y$ never crosses two sheets of the bifurcation set at a time. If this is not the case, perturb $y$ a little bit.
    Initially $SQD_{y}$ has three minima $m_{1}$, $m_{2}$, and $m_{3}$, lying on the faces incident to $V$ , three saddles $s_{1}$, $s_{2},s_{3}$, lying on edges that correspond to $A_{1},A_{2},A_{3}$, and one maximum $V$. Since $V$ is not the global maximum, there is at least one other maximum. Altogether there are at least 8 critical points of $SQD_{y}$, or, equivalently, 8 normals emanating from $y$.
@@
    (c) One of the saddles $s_{i}$ dies. This happens in two cases: (i) $s_{i}$ meets a maximum of $SQD_{y}$. The maximum cannot be $V$ since $s_{i}$ slides away from $V$. It is not any other maximum $V^{\prime}$. Indeed, recall that $s_{i}$ is the orthogonal projection of $y$ to the edge. If $s_{i}$ meets $V^{\prime}$ before $y$ leaves $\mathbf{P}$, then $|Vy|>|VV^{\prime}|$. Then $V^{\prime}$ can’t be maximum of $SQD_{y}$ when $y=V$. (ii) $s_{i}$ meets a minimum. It can be none of $m_{1}$, $m_{2}$, and $m_{3}$, so right before the bifurcation, the function $SQD_{y}$ has $4$ minima and $2$ maxima, which proves the claim.
@@
    Each skew polygon $P$ has exactly two acute angles, say, at vertices $M_{1}$ and $M_{2}$,
@@
Assume the contrary: a generic polytope $\mathbf{P}\subset\mathbb{R}^{3}$ has no point in its interior with 10 emanating normals. By Proposition 1, all the vertices of $\mathbf{P}$ are skew. Then, by Proposition 2 (1), each vertex of $\mathbf{P}$ is incident to exactly two acute edges. Therefore, acute edges form a closed broken line on the boundary of $\mathbf{P}$. Denote the broken line by $\mathcal{A}$.
@@
    $SQD_{y}:\partial\mathbf{P}\rightarrow\mathbb{R}.$
@@
(2) One way is clear: a maximum of $SQD_{y}$ is also a maximum of $SQD_{y,\mathcal{A}}$. The converse follows from Proposition 2, (3) and Lemma 5. ∎
@@
There exists a binormal $VW$ of $\mathbf{P}$ such that $V$ is a vertex, and $W$ lies in the interior of some edge.
@@
Binormals of $\mathbf{P}$ correspond to critical points of the width function of $\mathbf{P}$.  [^2] By definition, width function of $\mathbf{P}$ equals the distance between two parallel support planes to $\mathbf{P}$. The width function is clearly not a smooth function. So our aguments rely on the concept of Clarke subdifferential borrowed from non-smooth analysis, see e.g. [1], which contains also the definition of critical point and the regular interval theorem.
@@
(1) If the endpoints of a binormal are two vertices of $\mathbf{P}$, then the binormal realises a maximum of the width function.
(2) If the endpoints of a binormal are a vertex of $\mathbf{P}$ and a point in the interior of a face, then the binormal realises a minimum.
(3) If the endpoints of a binormal lie in the interiors of some (non-parallel, because of genericity) edges of $\mathbf{P}$, then the binormal also realises a minimum.
@@
Observe that $VW$ belongs to the active region $\mathcal{AR}(V)$, that is, for every point on the binormal, $V$ is a maximum.
@@
Take a point $y$ lying on the binormal very close to the vertex $V$. By Lemmata 5 and 3, $SQD_{y}$ has two saddles $s_{1},s_{2}$ on the acute edges incident to $V$, a saddle $s_{3}=W$, two minima $m_{1},m_{2}$ on the faces incident to $V$, and a maximum at $V$. Since $V$ is not a global maximum, there is some other global maximum; altogether there are at least $8$ normals emanating from $y$.
Figure 3. Binormal VWVW
@@
Let $VV^{\prime}$ and $VV^{\prime\prime}$ be two acute edges incident to $V$. For a point $y$ lying close to $V$ none of $V^{\prime}$ and $V^{\prime\prime}$ is a maximum of $SQD_{y}$.
@@
Keep moving until the first event, which is either a bifurcation (that is, a crossing of a sheet of $\mathcal{B}(\mathbf{P})$), or the point $y$ reaches $W$. We have the following case analysis.
@@
      One of the saddles $s_{i}$ meets a maximum, and the two critical points die. This saddle cannot be $s_{3}$ since $s_{3}$ does not move. It can be neither $s_{1}$ nor $s_{2}$ by Lemma 8.
@@
One repeats the arguments from Case 1 for a point $y$ travelling along the binormal shifted a little bit inwards the polytope. The same bifurcation arguments survive: if $y$ crosses a sheet of $\mathcal{B}(\mathbf{P})$ on its way, we get $10$ emanating normals.
So the only remaining case is the one when there are no bifurcations on the (shifted) binormal. Here we need additional arguments since one of the two new minima coincides with either $m_{1}$ or $m_{2}$. However we conclude that for a point $y$ on the (shifted) binormal we have (at least) $3$ minima, $2$ maxima, and $3$ saddles. Moreover, these $3$ saddles are attained on acute edges of $\mathbf{P}$. That is, we have three minima of $SQD_{y,\mathcal{A}}$, see Lemma 6. Since $\mathcal{A}$ is a topological circle (or a disjoint union of topological circles), there are necessarily three maxima of $SQD_{y,\mathcal{A}}$. By Lemma 6, they are maxima of $SQD_{y}$ as well.
@@ 4. Proof of Proposition 2
A bigon (or a lune) is a spherical polygon with two edges. It is bounded by two halves of big circles. The minimal bigon $Big(P)$ of a spherical polygon $P$ is the bigon containing $P$ which is minimal by inclusion, or, equivalently, with the minimal angle between its edges.
@@
Let $A$ be a vertex of $P$. Set $\mathcal{AR}^{spher}(A)$ be the set of points $Y\subset P$ such that $A$ is a short maximum of $SQD_{Y}^{spher}$.
@@
It remains to analyze triangles and quadrilaterals.
@@
[…]